# Explicit estimates for the logarithmic derivative and the reciprocal of the Riemann zeta-function

### v3

@@ 3 Some preliminaries
$t^{1/6}\log t\geq\frac{0.618(1+a_{0})(1+a_{1})^{1/6}}{\zeta(1+\delta_{r})}.$ (11)
@@
Recall that Hiary, Patel, and Yang [15] have $|\zeta(1/2+it)|\leq 0.618t^{1/6}\log t$, for $t\geq 3$.
We now refer to Lemma 4, taking the maximum of (9) at $\sigma=1/2$. It has been shown in [30, p. 15] that $Q_{0}=1.31$ is permissible in (8) together with
$(k_{1},k_{2},k_{3},Q_{0})=\left(0.618,\tfrac{1}{6},1,1.31\right),$
@@
We use same argument in the first paragraph of [29, p. 2], which modifies Ford’s proof in [10]. The only difference is that we now apply the argument to [2, §5], which itself is an improvement on Ford’s result. This proves the lemma. ∎
@@
with
$C_{3}=58.096\sqrt{1+9t_{0}^{-2}}\left(1+\frac{\log\sqrt{1+9t_{0}^{-2}}}{\log t_{0}}\right).$ (12)
@@
The proof is similar to that of [14, Lemma 7]. First, to show that the result holds for $1\leq\sigma\leq 2$, we use the Phragmén–Lindelöf Principle, i.e., Lemma 3, on the holomorphic function
@@
This inequality is verified numerically for $|t|<3$, and is a consequence of Lemma 5 for $|t|\geq 3$.
Also, on the line $s=2+it$,
@@
$\begin{split}\frac{|1+s|}{|s-1|}&\leq\frac{\sqrt{9+t^{2}}}{t}\leq\sqrt{1+9t_{0}^{-2}},\\ \log|1+s|&\leq\log\sqrt{9+t^{2}}\leq\log t+\log\sqrt{1+9t_{0}^{-2}}.\end{split}$
@@
$|\zeta(\sigma+it)|$ $\leq 58.096\sqrt{1+9t_{0}^{-2}}\left(1+\frac{\log\sqrt{1+9t_{0}^{-2}}}{\log t_{0}}\right)\log^{2/3}t<C_{3}\log^{2/3}t,$
@@
Finally, we address the case where $\sigma\geq 2$. Here since $\sigma>1$, we apply the trivial bound in (49). Then for $\sigma\geq 2$, it is easy to verify that
@@
[23, Lemma 4] In the region 1/$21/2<\sigma\leq 3/$ 2 and $0\leq t\leq 2\exp(e^{2})$, we have
@@
In the remainder of this section, we present some estimates concerning the Riemann zeta function, including some for $\sigma>1$, which are better than the trivial ones. We would like to point out that Lemma 8 and Corollary 9, which can both be viewed as an extension of Corollary 8 inside the critical strip, are not actually required for the purposes of this paper. However, we include them here as the proofs are similar to that of Corollary 8, and at the same time we provide a brief discussion on their usefulness. Lemma 8 is an improvement on [16, Proposition 5.3.2], which proves a similar result.
@@
where $\varphi^{*}(u)$ is the periodic function obtained by extending the polynomial $\varphi(u)=u^{2}-u+1/6$ on the interval $[0,1]$. Note that (14) holds for $\sigma>0$ by analytic continuation (see [26, Theorem 3.5]).
@@
This bound can be used throughout (14). The first term in (14) can also be bounded using the estimate in [11, Lemma 2.1], so for $N\geq 2$,
@@
$|\zeta(\sigma+ikt)|<\log kt+C(\sigma_{1},t_{0},k,\eta),$
@@
The proof follows the argument of [3, Proposition A.1], originally due to Backlund [1], and extends their result to $t<500$. We omit the proof since it is almost identical to that of Lemma 8. ∎
@@
Corollary 9 can be used in conjunction with other estimates for $\zeta(s)$ in the critical strip, such as Ford’s Theorem 1 in [10] or Trudgian’s Lemma 2.1 [28]. Ford’s bound, while asymptotically better, is only sharper than bounds of the form $\zeta(s)\ll\log t$ for very large $t$. For $\sigma\in[1-1/(W_{0}\log t),1]$ and $3\leq t\leq 10^{10^{5}}$, Corollary 9 is an improvement to [10, Theorem 1]. This makes it useful for results over smaller $t$.
@@
whereas degrading a Richert-type bound like Ford’s [10, Theorem 1] would give
@@
and a bound obtained via the Phragmén–Lindelöf principle like (2.4) in [28, Lem. 2.1] using estimates on the $1/2$-line (see Hiary, Patel, and Yang [15]) and estimates on the $1$-line (see Backlund [1]) would yield
@@
Next we have some bounds on $\zeta(\sigma)$ for real $\sigma>1$. These are useful in their own right, but we show they can be used to obtain good bounds on $-\zeta^{\prime}(\sigma)/\zeta(\sigma)$ in a similar region.
@@
[22, Lemma 5.4] Let $\gamma$ denote Euler’s constant. For $\sigma>1$ we have
@@
Let $\gamma_{n}$ be the Laurent–Stieltjes constants (with $\gamma:=\gamma_{0}$). For integers $k$ such that $1\leq k\leq 1500$, and real $\sigma$ satisfying $31<\sigma\leq\sigma_{k}<$ 3, we have
@@
Recall from the Laurent series expansion of the zeta function
$\zeta(\sigma)=\frac{1}{\sigma-1}+\gamma+\sum_{n=1}^{k}\frac{(-1)^{n}(\sigma-1)^{n}\gamma_{n}}{n!}+\sum_{n=k+1}^{\infty}\frac{(-1)^{n}(\sigma-1)^{n}\gamma_{n}}{n!}.$ (20)
To prove the non-negativity of (19) we require the last two terms of (20) together to be positive. We have the bounds on $\gamma_{n}$ due to Lavrik [17]:
@@
It is likely that $\phi(\sigma,k)$ is always positive for any $\sigma>1$, although proving that would require a different approach, perhaps one akin to [22, Lemma 5.3]. Our method in Lemmas 10 and 11, while simpler, is restricted by bounds on $|\gamma_{n}|$ which loses some information about sign changes. It is worth noting that bounds sharper than (21) are available, but are more complicated to work with.
@@
$-\zeta(\sigma)\leq\frac{1}{(\sigma-1)^{2}}+\frac{1}{(\sigma-3)^{2}},\qquad 1<\sigma<3,$
@@
for $|1-\sigma|<2$ by identification with the polylogarithm $\textup{Li}_{-1}(z)=z/(1-z)^{2}$. Dividing by the bound in the Lemma 10 proves the result. ∎
@@
The two factors on the right-hand side can be bounded with Lemmas 9 and 8, respectively, with the aid of the bound
@@ 4.1 Assuming a partial Riemann hypothesis
In this section, we assume $RH(T)$, and estimate $|\zeta^{\prime}(s)/\zeta(s)|$ in a right-half plane: $\sigma\geq 1/2+\epsilon$, $t_{0}\leq t\leq T$.
$\sigma_{0}=1+\epsilon,\qquad e+1/2\leq t_{0}\leq T_{0}\leq T-1/2,$
and $\epsilon\leq 1/2$ is a parameter to be determined later. We then extend their radii into the critical strip, so that Theorem 1 applied to $f(s):=-\zeta^{\prime}(s)/\zeta(s)$ gives us the desired estimates. Due to $RH(T)$, $\zeta(s)$ is analytic and non-zero in the discs.
Choosing a fixed $R^{\prime}<\epsilon$, we have by Corollary 6 that for $1/2+\epsilon\leq\sigma\leq 3/2+\epsilon$,
$\left|{\textup{Re}}\frac{\zeta^{\prime}}{\zeta}(\sigma+iT_{0})\right|\leq M:=\frac{4}{\pi R^{\prime}}\max_{0\leq\theta\leq 2\pi}|\log|\zeta(\sigma+iT_{0}+R^{\prime}e^{i\theta})||,$ (26)
which we estimate, by Lemma 7 with $\delta_{r}=1/2+\epsilon$ and $T_{0}\geq t_{0}\geq e+R^{\prime}$. In the region $1/2\leq\sigma\leq 3/2+2\epsilon$, we have
$|\zeta(s)|\leq A_{0}(T_{0}+R^{\prime})^{1/6}\log(T_{0}+R^{\prime})\leq A_{1}T_{0}^{1/6}\log T_{0},$ (27)
where
$A_{1}$ $=A_{0}\left(1+\frac{R^{\prime}}{t_{0}}\right)^{1/6}\left(1+\frac{\log\left(1+\frac{R^{\prime}}{t_{0}}\right)}{\log t_{0}}\right),$ (28)
with
$A_{0}$ $=0.618\left(1+a_{1}\left(\frac{3}{2}+2\epsilon,1.31,t_{0}-R^{\prime}\right)\right)^{7/6}$ (29)
×$\qquad\qquad\bm{\times}\left(1+a_{0}\left(\frac{3}{2}+2\epsilon,1.31,t_{0}-R^{\prime}\right)\right).$
Combining (27), (29), and (28) with (26) thus gives for $1/2+\epsilon\leq\sigma\leq 3/2+\epsilon$,
$M\leq\frac{4}{\pi R^{\prime}}\left(\frac{1}{6}\log T_{0}+\log\log T_{0}+\log A_{1}\right).$ (30)
Now to bound $|\zeta^{\prime}(s_{0})/\zeta(s_{0})|$, we can use the trivial bound for $\sigma>1$ [12, Lemma 70.1], which gives us
$\left|\frac{\zeta^{\prime}(s_{0})}{\zeta(s_{0})}\right|\leq-\frac{\zeta^{\prime}(\sigma_{0})}{\zeta(\sigma_{0})}<\frac{1}{\sigma_{0}-1}.$ (31)
However, when bounding $|\zeta^{\prime}(s_{0})/\zeta(s_{0})|$ in an interval away from the $1$-line, that is to say, where $\sigma_{0}\leq\sigma\leq 1.5$ with $\sigma_{0}$ not tending to $1$, we can do better than the trivial bound by using Corollary 10. We have for $\sigma_{0}\leq\sigma\leq 1.5$,
$\left|\frac{\zeta^{\prime}(s_{0})}{\zeta(s_{0})}\right|\leq-\frac{\zeta^{\prime}(\sigma_{0})}{\zeta(\sigma_{0})}<\frac{\phi_{2}(\sigma_{0})}{\sigma_{0}-1}.$ (32)
Up to this point, the above with Theorem 1, where $M$ is as in (30), and $R=1/2$, will give a result that holds for
$1+\epsilon-\frac{\alpha}{2}\leq\sigma\leq 1+\epsilon+\frac{\alpha}{2},$
and for a result on any larger $\sigma$, we simply use the trivial bound in (31).
Finally, putting everything together and making the substitution $T_{0}\mapsto t$, we prove our result.
@@
Assume $RH(T)$. Let $\alpha$, $\epsilon$, and $R^{\prime}$ be parameters, such that $20<R^{\prime}<\epsilon\leq 1/$ 2 and $10<\alpha<$ 1. In the region
$\sigma\geq 1+\epsilon-\frac{\alpha}{2}\quad\text{and}\quad R^{\prime}+e\leq t_{0}\leq t\leq T-\frac{1}{2},$
@@
$Q_{RH(T)}$ $=\max\Bigg{\{}\frac{2\alpha}{(1-\alpha)\pi R^{\prime}}\Bigg{(}\frac{1}{6}+B_{t_{0}}+\frac{\log A_{1}}{\log t_{0}}\Bigg{)}$ (33)
$\qquad\qquad+\left(\frac{1+\alpha}{1-\alpha}\right)\frac{\phi_{2}(1+\epsilon)}{\epsilon\log t_{0}},\,\frac{1}{(\epsilon+(\alpha/2))\log t_{0}}\Bigg{\}},$
where $A_{1}$ is defined in (28), $\phi_{2}(\sigma)$ is defined in Corollary 10, and
$B_{t_{0}}=\begin{cases}\frac{1}{e}&\text{ if }t_{0}\leq e^{e},\\ \frac{\log\log t_{0}}{\log t_{0}}&\text{ if }t_{0}>e^{e}.\end{cases}$
@@
For $t\geq t_{0}$, in Theorem 2 set $\sigma_{0}=1+\epsilon-\frac{\alpha}{2}$. Then fix $t_{0}$ and optimise (33) over $\epsilon,\,R^{\prime},\,\alpha$, while ensuring that the restrictions set in Theorem 2 and (11) hold. In practice $R^{\prime}$ is taken to be arbitrarily smaller than $\epsilon$, differing only after the sixth decimal place. The final value of $t_{0}=5$ is taken to be the one where (11) holds for all $\sigma_{0}$ in Table 2. This gives the result. ∎
$\sigma_{0}$ $Q_{RH(H_{0})}$ $\epsilon$ $\alpha$
$0.51$ $38889$ $0.004948$ $0.989897$
$0.55$ $1508$ $0.024334$ $0.948668$
$0.6$ $361.726$ $0.047605$ $0.89521$
$0.65$ $153.957$ $0.070599$ $0.841198$
$0.7$ $82.739$ $0.092279$ $0.784558$
$0.75$ $50.441$ $0.113238$ $0.726476$
$0.8$ $33.249$ $0.129407$ $0.658814$
$0.85$ $23.086$ $0.144234$ $0.588468$
$0.9$ $16.601$ $0.160246$ $0.520492$
$0.95$ $12.235$ $0.174136$ $0.448272$
$1$ $9.161$ $0.176533$ $0.353067$
Table 2: Values for $Q_{RH(H_{0})}$ and $\sigma\geq\sigma_{0}$ in Corollary 3, with corresponding $\epsilon,\,\alpha$ (all rounded to six decimal places). Each entry is valid for $5\leq t\leq H$.

### v4

@@ 3 Some preliminaries
$t^{1/6}\log t\geq\frac{\zeta(1+\delta_{r})}{0.618(1+a_{1}(1+\delta_{r},1.31,t_{0}))^{1/6}},$ (6)
where $\delta_{r}$, $a_{0}(\sigma,1.31,t)$, and $a_{1}(\sigma,1.31,t)$ are defined as in Lemma 4.
@@
Recall that Hiary, Patel, and Yang [16] have $|\zeta(1/2+it)|\leq 0.618t^{1/6}\log t$, for $t\geq 3$, while a quick calculation on Mathematica shows that $|\zeta(1/2+it)|\leq 1.461$ for $|t|<3$ (see also [14]). Meanwhile, one has by $|\zeta(s)|\leq\zeta(\sigma)$ that for all $t$,
$|\delta_{r}+it||\zeta(1+\delta_{r}+it)|\leq\frac{\zeta(1+\delta_{r})}{\log t}|\delta_{r}+it|\log t.$
Therefore, it can be verified that both bounds in (3) are met.
We now refer to Lemma 4, taking the maximum of (4) at $\sigma=1/2$. It has been shown in [31, p. 15] that $Q_{0}=1.31$ is permissible in (3) together with
$(k_{1},k_{2},k_{3},k_{4},Q_{0})=\left(0.618,\tfrac{1}{6},1,\tfrac{\zeta(1+\delta_{r})}{\log t},1.31\right),$
@@
We use same argument in the first paragraph of [28, p. 260], which modifies Ford’s proof in [11, Lemma 7.3]. The only difference is that we now apply the argument with parameters $C$ and $D$ found in the proof of Theorem $1.1$ of [2, §5], which itself is an improvement on Ford’s result. This proves the lemma. ∎
@@
with $a_{0}$ defined as in Lemma 4, and
$C_{3}=58.096\sqrt{1+9t_{0}^{-2}}\left(1+a_{0}(2,1,t_{0})\right)^{2/3}.$ (7)
@@
The proof is similar to that of [15, Lemma 7]. First, to show that the result holds for $1\leq\sigma\leq 2$, we use the Phragmén–Lindelöf Principle, i.e., Lemma 3, on the holomorphic function
@@
This inequality is verified numerically via Mathematica for $|t|<3$, and is a consequence of Lemma 5 for $|t|\geq 3$.
Also, on the line $s=2+it$, the trivial bound $|\zeta(s)|\leq\zeta(\sigma)$ gives
@@
$\begin{split}\frac{|1+s|}{|s-1|}&\leq\frac{\sqrt{9+t^{2}}}{t}\leq\sqrt{1+9t_{0}^{-2}},\\ |\log(1+s)|&\leq\left(1+a_{0}(2,1,t_{0})\right)\log t,\end{split}$
where $a_{0}$ is defined as in Lemma 4. The second inequality above is derived using the same argument as that in [30, §2].
@@
$|\zeta(\sigma+it)|$ $\leq 58.096\sqrt{1+9t_{0}^{-2}}\left(1+a_{0}(2,1,t_{0})\right)^{2/3}\log^{2/3}t\leq C_{3}\log^{2/3}t,$
@@
Finally, we address the case where $\sigma\geq 2$. Here since $\sigma>1$, we apply the trivial bound
$|\zeta(s)|\leq\zeta(\sigma)\leq\frac{\sigma}{\sigma-1},$ (9)
where $\sigma/(\sigma-1)$ is decreasing in $\sigma$. Then for $\sigma\geq 2$, it is easy to verify that
@@
[24, Lemma 4] In the region $1/2<\sigma\leq 3/2$ and $0\leq t\leq 2\exp(e^{2})$, we have
@@
In the remainder of this section, we present some estimates concerning the Riemann zeta function, including some for $\sigma>1$, which are better than the trivial ones in some finite range.
We would like to point out that Lemmas 10 and 11, as well as Corollary 10, do not contribute to the main results of this paper, but are included as they are potentially useful estimates for anyone seeking to do explicit computations involving $\zeta(s)$.
Similarly, Lemma 8 and Corollary 9, which can both be viewed as an extension of Corollary 8 inside the critical strip, are not actually required for the main purposes of this paper. However, we include them here as the proofs are similar to that of Corollary 8, and at the same time we provide a brief discussion on their usefulness. Lemma 8 is an improvement on [17, Proposition 5.3.2], which proves a simlar result.
@@
where $\varphi^{*}(u)$ is the periodic function obtained by extending the polynomial $\varphi(u)=u^{2}-u+1/6$ on the interval $[0,1]$ (for full details see e.g., [1, Equation (8)]). Note that (10) holds for $\sigma>0$ by analytic continuation (see [26, Theorem 3.5]).
@@
This bound can be used throughout (10). The first term in (10) can also be bounded using the estimate in [12, Lemma 2.1], so for $N\geq 2$,
@@
$|\zeta(\sigma+ikt)|<\log(kt)+C(\sigma_{1},t_{0},k,\eta),$
@@
The proof follows the argument of [3, Proposition A.1], originally due to Backlund [1], and extends their result to $t<500$. Note that here we have instead the choice of $N=\lfloor kt/\eta\rfloor+1$. A slightly better estimate of the harmonic sum from [12, Lemma 2.1] was also used. We omit the proof since the process is almost identical to that of Lemma 8. ∎
@@
Corollary 9 can be used in conjunction with other estimates for $\zeta(s)$ in the critical strip, such as Ford’s Theorem 1 in [11] or Trudgian’s Lemma 2.1 [30]. Ford’s bound, while asymptotically better, is only sharper than bounds of the form $\zeta(s)\ll\log t$ for very large $t$. For $\sigma\in[1-1/(W_{0}\log t),1]$ and $3\leq t\leq 10^{10^{5}}$, Corollary 9 is an improvement to [11, Theorem 1]. This makes it useful for results over smaller $t$.
@@
whereas degrading a Richert-type bound like Ford’s [11, Theorem 1] would give
@@
and a bound obtained via the Phragmén–Lindelöf principle like (2.4) in [30, Lem. 2.1] using estimates on the $1/2$-line (see Hiary, Patel, and Yang [16]) and estimates on the $1$-line (see Backlund [1]) would yield
@@
Next we have some bounds on $\zeta(\sigma)$ for real $\sigma>1$. These are useful in their own right, but we show they can be used to obtain good bounds on $\zeta^{\prime}(\sigma)/\zeta(\sigma)$ in a similar region.
@@
[23, Lemma 5.4] Let $\gamma$ denote Euler’s constant. For $\sigma>1$ we have
@@
Let $\gamma_{n}$ be the Laurent–Stieltjes constants (with $\gamma:=\gamma_{0}$). For integers $k$ such that $1\leq k\leq 1500$, and real $\sigma$ satisfying $1<\sigma\leq\sigma_{k}<3$, we have
@@
To prove the first assertion of the lemma, we have from the Laurent series expansion (about $\sigma=1$) of the zeta function
$\zeta(\sigma)$ $=\frac{1}{\sigma-1}+\gamma+\sum_{n=1}^{k}\frac{(-1)^{n}(\sigma-1)^{n}\gamma_{n}}{n!}+\sum_{n=k+1}^{\infty}\frac{(-1)^{n}(\sigma-1)^{n}\gamma_{n}}{n!}$ (16)
$\geq\frac{1}{\sigma-1}+\gamma+\sum_{n=1}^{k}\frac{(-1)^{n}(\sigma-1)^{n}\gamma_{n}}{n!}-\left|\sum_{n=k+1}^{\infty}\frac{(-1)^{n}(\sigma-1)^{n}\gamma_{n}}{n!}\right|.$
We have the bounds on $\gamma_{n}$ due to Lavrik [18]:
@@
It is likely that $\phi_{0}(\sigma,k)$ is always positive for any $\sigma>1$, although proving that would require a different approach, perhaps one akin to [23, Lemma 5.3]. Our method in Lemmas 10 and 11, while simpler, is restricted by bounds on $|\gamma_{n}|$ which loses some information about sign changes. It is worth noting that bounds sharper than (17) are available, but are more complicated to work with.
@@
$\zeta^{\prime}(\sigma)\leq\frac{1}{(\sigma-1)^{2}}+\frac{1}{(\sigma-3)^{2}},\qquad 1<\sigma<3,$
@@
for $|1-\sigma|<2$ by identification with the polylogarithm $\textup{Li}_{-1}(z)=z/(1-z)^{2}$. Dividing by the bound in Lemma 10 proves the result. ∎
@@
The two factors on the right-hand side can be bounded with Lemma 9 and Corollary 8, respectively, with the aid of the bound
@@ 4.1 Assuming a partial Riemann hypothesis
In this section, we assume $RH(T)$, and estimate $|\zeta^{\prime}(s)/\zeta(s)|$ in a right-half plane: $\sigma>1/2+\epsilon$, $t_{0}\leq t\leq T$.
@@
Assume $RH(T)$. Let $\alpha_{0}$ and $\epsilon$ be parameters, such that $0<\epsilon\leq 1/2$ and $0<\alpha_{0}<1$. In the region
$\sigma\geq 1+\epsilon-\frac{\alpha_{0}}{2}\quad\text{and}\quad 3\leq t_{0}\leq t\leq T-\frac{1}{2},$
@@
$Q_{RH(T)}$ $=\max\Bigg{\{}\frac{4}{(1-\alpha_{0})^{2}}\Bigg{(}\frac{1}{6}+2B_{t_{0}}+\frac{\log A_{3}}{\log t_{0}}\Bigg{)}\,,\,\frac{1}{(\epsilon+(\alpha_{0}/2))\log t_{0}}\Bigg{\}},$ (22)
where $A_{3}$ is defined in (4.1), and
$B_{t_{0}}=\begin{cases}\frac{1}{e}&\text{ if }t_{0}\leq e^{e},\\ \frac{\log\log t_{0}}{\log t_{0}}&\text{ if }t_{0}>e^{e},\end{cases}$ (23)
provided that
$t^{1/6}\log t\geq\frac{\zeta(\frac{3}{2}+\epsilon)}{0.618(1+a_{1}(\frac{3}{2}+\epsilon,1.31,t_{0}))^{1/6}}$ (24)
is satisfied, with $a_{1}(\sigma,1.31,t)$ defined as in Lemma 4.
@@
Proof.
$\sigma_{0}=1+\epsilon,\qquad 3\leq t_{0}\leq T_{0}\leq T-1/2,$
and $\epsilon\leq 1/2$ is a parameter to be determined later. We then extend their radii into the critical strip, so that Corollary 6 applied with
$f(s):=\zeta(s),\qquad s_{0}=\sigma_{0}+iT_{0},\qquad r=1/2,$
gives us the desired estimates. Due to $RH(T)$, $\zeta(s)$ is analytic and non-zero in the discs.
In order to do this, we first need to determine $A_{1}$. Let $V(\epsilon,\sigma_{1},t_{0},\eta)$ be defined as in Lemma 12, with $\sigma_{1}$ and $\eta$ as parameters. Then by Corollary 7 with $\delta_{r}=r+\epsilon$, and Lemma 12 with $\kappa=\epsilon$, we have for $1/2\leq\sigma\leq\sigma_{0}+r$ and each $T_{0}\in[T_{0}-r,T_{0}+r]$,
$\left|\frac{\zeta(\sigma+iT_{0})}{\zeta(\sigma_{0}+iT_{0})}\right|\leq A_{1}:=A_{3}T_{0}^{1/6}(\log T_{0})^{2}$ (25)
where
$A_{3}:=0.618$ $(1+a_{0}(\sigma_{0}+r,1.31,t_{0}))(1+a_{1}(\sigma_{0}+r,1.31,t_{0}))^{7/6}$
$\bm{\cdot}\left(1+\frac{r}{t_{0}}\right)^{1/6}\left(1+\frac{\log\left(1+\frac{r}{t_{0}}\right)}{\log t_{0}}\right)V(\epsilon,\sigma_{1},t_{0},\eta);$ (26)
provided (6) is satisfied.
Now choosing a fixed positive $\alpha_{0}<1$, we have by Corollary 6,
$\left|\frac{\zeta^{\prime}(\sigma+iT_{0})}{\zeta(\sigma+iT_{0})}\right|$ $\leq\frac{4}{(1-\alpha_{0})^{2}}\left(\frac{1}{6}\log T_{0}+2\log\log T_{0}+\log A_{3}\right)$
$\leq\frac{4}{(1-\alpha_{0})^{2}}\left(\frac{1}{6}+2B_{t_{0}}+\frac{\log A_{3}}{\log t_{0}}\right)\log T_{0},$ (27)
which holds for
$1+\epsilon-\frac{\alpha_{0}}{2}\leq\sigma\leq 1+\epsilon+\frac{\alpha_{0}}{2}.$
Note that $B_{t_{0}}$ is defined in (23), and is utilised since $\log\log t/\log t$ attains its maximum at $t=e^{e}$, after which it is decreasing.
To obtain a bound that is valid also for $\sigma\geq 1+\epsilon+\alpha_{0}/2$, we use the trivial bound for all $\sigma>1$ [13, Lemma 70.1]:
$\left|\frac{\zeta^{\prime}(s)}{\zeta(s)}\right|\leq-\frac{\zeta^{\prime}(\sigma)}{\zeta(\sigma)}<\frac{1}{\sigma-1}$ (28)
with $\sigma=1+\epsilon+\alpha_{0}/2$, and then take the maximum of (4.1) and (28).
Finally, making the substitution from $T_{0}$ to $t$ proves the result. ∎
@@
For $t\geq t_{0}$ and $\sigma\geq\sigma_{0}$, in Theorem 2 set $\sigma_{0}=1+\epsilon-\frac{\alpha_{0}}{2}$. Then fixing $t_{0}$ and $\sigma_{0}$, optimise (22) over $\epsilon$ (which also determines $\alpha_{0}$), and also over $\sigma_{1},\,\eta$ (due to the quantity $V$ in $A_{3}$), while ensuring that the restrictions set in Theorem 2 (and by extension Lemma 12 and Corollary 8), hold. The value $t_{0}$ is chosen by finding the smallest value of $t$ that satisfies (24). ∎
$\sigma_{0}$ $t_{0}$ $Q_{RH(H_{0})}$ $\epsilon$ $\sigma_{1}$ $\eta$
$0.51$ $15$ $31447$ $0.000462$ $1.965754$ $2.608856$
$0.55$ $14$ $1086$ $0.002766$ $1.901324$ $3.329589$
$0.6$ $14$ $249.260$ $0.006041$ $1.795116$ $3.162609$
$0.65$ $14$ $104.992$ $0.009573$ $1.846979$ $3.195987$
$0.7$ $14$ $56.739$ $0.013287$ $1.797626$ $3.198866$
$0.75$ $14$ $35.162$ $0.017144$ $1.155376$ $2.868261$
$0.8$ $14$ $23.759$ $0.021126$ $1.392644$ $3.173843$
$0.85$ $14$ $17.050$ $0.025202$ $1.763188$ $3.402204$
$0.9$ $13$ $13.123$ $0.029430$ $1.227811$ $3.283187$
$0.95$ $13$ $10.175$ $0.033682$ $1.227876$ $3.283802$
11 $13$ $8.101$ $0.037999$ $1.889284$ $3.054339$
Table 2: Values for $Q_{RH(H_{0})}$ and $\sigma\geq\sigma_{0}$ in Corollary 3, with corresponding $\epsilon,\,\sigma_{1},\,\eta$ (all rounded to six decimal places). Each entry is valid for $t_{0}\leq t\leq H$.
We note that [5, Theorem 1.1] implies that
$\left|\frac{\zeta^{\prime}(1+it)}{\zeta(1+it)}\right|\leq 0.639\log t$
for $10^{6}\leq t\leq 2.99997\cdot 10^{12}$. This result allows us to show that
$\left|\frac{\zeta^{\prime}(1+it)}{\zeta(1+it)}\right|\leq 40.440\log\log t$ (29)
holds unconditionally for $e^{e}\leq t\leq 2.99997\cdot 10^{12}$.
To prove this, recall from §2.1 that $RH$ has been verified up to $3.000175332800\cdot 10^{12}$. Now note that $\log t/\log\log t$ is increasing for $t\geq e^{e}$. Then repeating the process in the proof of Corollary 3, for $e^{e}\leq t\leq 10^{6}$, we obtain
$\left|\frac{\zeta^{\prime}(1+it)}{\zeta(1+it)}\right|\leq 7.686\log t\leq 40.440\log\log t,$
with rounded parameters $(\epsilon,\sigma_{1},\eta)=(0.037816,1.395842,3.177141)$. The second inequality above uses the fact that $t$ attains its maximum at $10^{6}$.
We remark that (29) is an unconditional result due to the partial verification of $RH$. While this is nowhere near as sharp as the bounds in [5], it nevertheless holds for a wider range and might be useful for smaller values of $t$.
If desired, one can extend the range for which the corollary is unconditional beyond this (albeit still in a finite range $t_{0}\leq t\leq T$), at the expense of a worse leading constant. This will prove a poor man’s version of [5, Theorem 1.1], but with a simpler method. As mentioned in §1, the range of validity is wider but the constants are worse. This extension requires a modification in the proof of Theorem 2, by fixing $r=\epsilon+1/W\log T$, utilising a classical zero-free region, and also using Lemma 8 in place of Corollary 7 when estimating $A_{1}$ in (25).